sam (.sam.):

Prove (cos(x)+2sin(x))^2 + (2cos(x)-sin^2(x))^2 = 5 Not sure if provable or not.

OpenStudy (rogue):

According to wolfram, nope.

OpenStudy (rogue):

When x = 0, the thing does = 5, but simplifying the left side down to 5 is not possible.

sam (.sam.):

yeah, thought of that, and this equation makes me solve for half a page, lol

OpenStudy (aravindg):

help pls

OpenStudy (zarkon):

did you mean \[(\cos(x)+2\sin(x))^2 + (2\cos(x)-\sin(x))^2 = 5\]